# Sobolev regularity theory for stochastic reaction-diffusion-advection equations with spatially homogeneous colored noises and variable-order nonlocal operators

### v2

@@ Abstract.
This article investigates the existence, uniqueness, and regularity of solutions to nonlinear stochastic reaction-diffusion-advection equations (SRDAEs) with spatially homogeneous colored noises and variable-order nonlocal operators in mixed norm $L_{q}(L_{p})$-spaces. We introduce a new condition—strongly reinforced Dalang’s condition—on colored noise, which facilitates a deeper understanding of the complicated relation between nonlinearities and stochastic forces. Additionally, we establish the space-time Hölder type regularity of solutions.
@@
Stochastic reaction-diffusion-advection equation, Colored noise, Variable-order nonlocal operator, Sobolev regularity, Mixed norm, Hölder regulairty
@@ 1.1. Goal and Settings
Here $\overline{z}$ is a complex conjugate of $z$.
@@ 1.2. Background
When $d=1$, $\mathcal{L}$ is a second-order elliptic operator, and $\dot{W}$ is the space-time white noise, C. Mueller [44] established that for $0\leq\lambda_{\text{s.m.}}<\frac{1}{2}$, a mild solution exists globally in time on a finite spatial interval with Dirichlet boundary conditions. Moreover, C. Mueller [45] also demonstrated that for $\lambda_{\text{s.m.}}>\frac{1}{2}$, the mild solution exhibits finite-time explosion with non-zero probability under the same conditions. In the critical scenario where $\lambda_{\text{s.m.}}=\frac{1}{2}$, M. Salin [53] showed that solution explosions are precluded within a periodic domain. In instances where non-explosive solutions are considered, for $0\leq\lambda_{\text{s.m.}}<\frac{1}{2}$, significant results have emerged. Notably, N.V. Krylov [39, Section 8.4] established the existence and Sobolev regularity of the strong solution over $\mathbb{R}$. B.-S. Han [26] confirmed the existence, uniqueness, Sobolev regularity, and boundary behavior of strong solutions on a finite interval with Dirichlet boundary conditions. Additionally, B.-S. Han [27] further substantiated the existence, uniqueness, and Sobolev regularity of strong solutions over $\mathbb{R}$ with $\mathcal{L}=a\Delta^{\alpha/2}+bD+c$. These results predominantly address the dynamics influenced by space-time white noise. J.-H. Choi and B.-S. Han [9] demonstrated the existence, uniqueness, and Sobolev regularity of strong solutions in $\mathbb{R}^{d}$ with spatially homogeneous colored noise $\dot{W}$ and $\mathcal{L}=a^{ij}D^{ij}+b^{i}D^{i}+c$. For further discussions on the range $-1<\lambda_{\text{s.m.}}<0$, the reader is referred to [4, 46, 47].
@@
Assuming $\mathcal{L}$ as a second-order elliptic operator and $\dot{W}$ as spatially homogeneous colored noise, the scenario where $\varphi$ exhibits linear growth and $F$ has polynomial growth with a negative leading coefficient has been addressed by S. Cerrai, G. Da Prato, and F. Flandoli [5, 6, 7]. They proved the existence and uniqueness of mild solutions over bounded open sets in $\mathbb{R}^{d}$ with regular boundaries. Additionally, in the case where $|F(u)|\lesssim|u|\log|u|$ and $|\varphi(u)|\lesssim|u|(\log|u|)^{\kappa/2}$ for some $\kappa>0$ and large $|u|$, R.C. Dalang, D. Khoshnevisan, and T. Zhang [16] established the existence and uniqueness of the mild solution with space-time white noise. L. Chen and J. Huang [8] extended these works to settings with spatially homogeneous colored noise under reinforced Dalang’s condition. When $F(u)\text{sign}(u)\lesssim-|u|^{1+\beta}$ and $|\varphi(u)|\lesssim(1+|u|)^{1+\gamma}$, M. Salin [52] confirmed the existence of a global mild solution. Inspired by Salin’s findings, which suggest that larger values of $\gamma$ can be chosen if $\beta$ is substantial, B.-S. Han and J. Yi [30] validated the existence, uniqueness, and Sobolev regularity of strong solutions with $F(u)=-|u|^{1+\beta}$ and $\varphi(u)=|u|^{1+\gamma}$ in contexts with spatially homogeneous colored noise meeting reinforced Dalang’s condition. Recently, A. Agresti and M.C. Veraar [1, 2] demonstrated the existence, uniqueness, and regularity of solutions within critical spaces for stochastic reaction-diffusion equations with transport noise in periodic domains.
@@
    Non-local operators: Beyond the findings of B.-S. Han [27], research has predominantly focused on second-order elliptic operators. This observation prompts the question: Is it feasible to establish solvability for SRDAEs that involve variable-order non-local operators?
@@
This article is dedicated to addressing the above questions, aiming to fill the gaps identified in current research on SRDAEs.
@@ 1.3. Conditions on spatially homogeneous colored noise
This subsection details the conditions imposed on spatially homogeneous colored noise $W$ characterized by a nonnegative definite correlation $\pi$. Here is our main assumption on $\pi$. Throughout this article, we always set
$\delta_{0}\in(0,1],\quad\gamma\in(0,1)\quad\text{and}\quad r\in[1,d/(d-\delta_{0}(1-\gamma)).$
@@
    The corresponding measure $\mu$ of $\pi$ (see [9, Theorem 2.2]) satisfies that
    $\begin{cases}\int_{|x|<1}|x|^{2\alpha_{\delta_{0},\gamma,r}-d}\mu(\mathrm{d}x)<\infty,\quad&\text{if }0<\alpha_{\delta_{0},\gamma,r}<d/2,\\ \int_{|x|<1}\log(1/|x|)\mu(\mathrm{d}x)<\infty,\quad&\text{if }\alpha_{\delta_{0},\gamma,r}=d/2,\\ \text{no specific conditions on $\mu$,}\quad&\text{if }\alpha_{\delta_{0},\gamma,r}>d/2,\end{cases}$
@@
$\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}\ast R_{\delta_{0}(1-\gamma)})(x)\mu(\mathrm{d}x)=\int_{\mathbb{R}^{d}}R_{2\delta_{0}(1-\gamma)}(x)\mu(\mathrm{d}x)=:\nu_{\delta_{0},\gamma,1}<\infty,$
@@
Let $\delta_{0}\in(0,1]$, $\gamma\in(0,1)$, $r_{1},r_{2}\in[1,d/(d-\delta_{0}(1-\gamma))$ and $\mu$ be a tempered measure defined on $\mathbb{R}^{d}$. Suppose that $r_{1}<r_{2}$. Then, Assumption 1.5 $(\delta_{0},\gamma,r_{2})$ implies Assumption 1.5 $(\delta_{0},\gamma,r_{1})$.
@@
    for all $x\in\mathbb{R}^{d}\setminus\{0\}$,
    $R_{\delta_{0}(1-\gamma)}(x)\leq N(d,\delta_{0},\gamma)|x|^{\delta_{0}(1-\gamma)-d}1_{|x|\leq 2}+N(d,\delta_{0},\gamma)\mathrm{e}^{-\frac{|x|}{2}}1_{|x|>2}.$ (1.5)
@@
    for $|x|\leq 2$,
    $N(d,\delta_{0},\gamma)|x|^{\delta_{0}(1-\gamma)-d}\leq R_{\delta_{0}(1-\gamma)}(x).$
@@
$(4^{\prime})$ for $\kappa>0$, there exists $N=N(d,\delta_{0},\gamma,\kappa)>$ such that for $|x|<\kappa$,
$N|x|^{\delta_{0}(1-\gamma)-d}\leq R_{\delta_{0}(1-\gamma)}(x).$ (1.6)
@@
Let $\delta_{0}\in(0,1]$, $\gamma\in(0,1)$, $r\in[1,d/(d-\delta_{0}(1-\gamma))$ and $\mu$ be a tempered measure defined on $\mathbb{R}^{d}$. Then, the following conditions are equivalent:
@@
    The convolution $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})$ belongs to $L_{1}(\mathbb{R}^{d},\mu)$;
    $\nu_{\delta_{0},\gamma,r}:=\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)\mu(\mathrm{d}x)<\infty.$ (1.8)
@@
    The convolution $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})$ belongs to $L_{1}(B_{1}(0),\mu)$;
    $\int_{|x|<1}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)\mu(\mathrm{d}x)<\infty.$ (1.9)
@@
The implication from condition $(ii)$ to $(iii)$ is straightforward. By examining the exponential decay of $R_{\delta_{0}(1-\gamma)}$ as $|x|\to\infty$, as outlined in (1.5) and corroborated by [9, Lemma 4.3], we infer that the convolution $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)$ exhibits similar exponential decay for large $x$. Furthermore, the characterization of $\mu$ as a tempered measure, detailed in [9, Definition 2.1], stipulates the existence of some $k\geq 0$ such that
$A_{\mu}:=\int_{\mathbb{R}^{d}}\frac{1}{(1+|x|^{2})^{k/2}}\mu(\mathrm{d}x)<\infty.$ (1.10)
This integral condition on $\mu$ robustly supports the assertion that $(iii)$ implies $(ii)$. We proceed by establishing the equivalence between conditions $(i)$ and $(iii)$ through three cases based on the value of $\alpha_{\delta_{0},\gamma,r}$.
@@
where the strict positivity and radial decrease of $R_{\delta_{0}(1-\gamma)}$ imply for $31/2<|y|<$ 3,
@@
$\nu_{\delta_{0},\gamma,r_{1}}$ $:=\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r_{1}}\ast R_{\delta_{0}(1-\gamma)}^{r_{1}})(x)\mu(\mathrm{d}x)$ (1.13)
$\leq C_{0}\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\mu(\mathrm{d}x).$
Given that $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)$ decays exponentially as $|x|\to\infty$ and considering $\mu$ as a tempered measure,
$\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\mu(\mathrm{d}x)<\infty\Longleftrightarrow\int_{|x|\leq 1}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\mu(\mathrm{d}x)<\infty.$ (1.14)
@@
$\int_{|x|\leq 1}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\mu(\mathrm{d}x)$ (1.15)
$\leq A_{\mu}^{1-\frac{r_{1}^{2}}{r_{2}^{2}}}\left(\int_{|x|\leq 1}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)(1+|x|^{2})^{\frac{k}{2}\left(\frac{r_{2}^{2}}{r_{1}^{2}}-1\right)}\mu(\mathrm{d}x)\right)^{\frac{r_{1}^{2}}{r_{2}^{2}}}$
$\leq N(A_{\mu},k,r_{1},r_{2})\nu_{\delta_{,}\gamma,r_{2}}^{\frac{r_{1}^{2}}{r_{2}^{2}}},$
where $A_{\mu}$ is the constant defined in (1.10). Consequently,

### v3

# Sobolev regularity theory for stochastic reaction-diffusion-advection equations with spatially homogeneous colored noises and infinitesimal generators of subordinate Brownian motions

@@ Abstract.
This article investigates the existence, uniqueness, and regularity of solutions to nonlinear stochastic reaction-diffusion-advection equations (SRDAEs) with spatially homogeneous colored noises and infinitesimal generators of subordinate Brownian motions in mixed norm $L_{q}(L_{p})$-spaces. We introduce a new condition—strongly reinforced Dalang’s condition—on colored noise, which facilitates a deeper understanding of the complicated relation between nonlinearities and stochastic forces. Additionally, we establish the space-time Hölder type regularity of solutions.
@@
Stochastic reaction-diffusion-advection equation, Colored noise, Infinitesimal generator, Sobolev regularity, Mixed norm, Hölder regularity
@@ 1.1. Goal and Settings
Here $\overline{z}$ is the complex conjugate of $z$.
@@ 1.2. Background
When $d=1$, $\mathcal{L}$ is a second-order elliptic operator, and $\dot{W}$ is the space-time white noise, C. Mueller [44] established that for $0\leq\lambda_{\text{s.m.}}<\frac{1}{2}$, a mild solution exists globally in time on a finite spatial interval with periodic boundary conditions. Moreover, C. Mueller [45] also demonstrated that for $\lambda_{\text{s.m.}}>\frac{1}{2}$, the mild solution exhibits finite-time explosion with non-zero probability under the same conditions. In the critical scenario where $\lambda_{\text{s.m.}}=\frac{1}{2}$, M. Salins [54] showed that solution explosions are precluded within a periodic domain. In instances where non-explosive solutions are considered, for $0\leq\lambda_{\text{s.m.}}<\frac{1}{2}$, significant results have emerged. Notably, N.V. Krylov [39, Section 8.4] established the existence and Sobolev regularity of the strong solution over $\mathbb{R}$. B.-S. Han [26] confirmed the existence, uniqueness, Sobolev regularity, and boundary behavior of strong solutions on a finite interval with Dirichlet boundary conditions. Additionally, B.-S. Han [27] further substantiated the existence, uniqueness, and Sobolev regularity of strong solutions over $\mathbb{R}$ with $\mathcal{L}=a\Delta^{\alpha/2}+bD+c$. These results predominantly address the dynamics influenced by space-time white noise. J.-H. Choi and B.-S. Han [9] demonstrated the existence, uniqueness, and Sobolev regularity of strong solutions in $\mathbb{R}^{d}$ with spatially homogeneous colored noise $\dot{W}$ and $\mathcal{L}=a^{ij}D^{ij}+b^{i}D^{i}+c$. For further discussions on the range $1<\lambda_{\text{s.m.}}<0$, the reader is referred to [4, 46, 47].
@@
Assuming $\mathcal{L}$ as a second-order elliptic operator and $\dot{W}$ as spatially homogeneous colored noise, the scenario where $\varphi$ exhibits linear growth and $F$ has polynomial growth with a negative leading coefficient has been addressed by S. Cerrai, G. Da Prato, and F. Flandoli [5, 6, 7]. They proved the existence and uniqueness of mild solutions over bounded open sets in $\mathbb{R}^{d}$ with regular boundaries. Additionally, in the case where $|F(u)|\lesssim|u|\log|u|$ and $|\varphi(u)|\lesssim|u|(\log|u|)^{\kappa/2}$ for some $\kappa>0$ and large $|u|$, R.C. Dalang, D. Khoshnevisan, and T. Zhang [16] established the existence and uniqueness of the mild solution with space-time white noise. L. Chen and J. Huang [8] extended these works to settings with spatially homogeneous colored noise under reinforced Dalang’s condition. When $F(u)\text{sign}(u)\lesssim-|u|^{1+\beta}$ and $|\varphi(u)|\lesssim(1+|u|)^{1+\gamma}$, M. Salins [53] confirmed the existence of a global mild solution. Inspired by the result of Salins [53], which suggests that larger values of $\gamma$ can be chosen if $\beta$ is substantial, B.-S. Han and J. Yi [30] validated the existence, uniqueness, and Sobolev regularity of strong solutions with $F(u)=-|u|^{1+\beta}$ and $\varphi(u)=|u|^{1+\gamma}$ in contexts with spatially homogeneous colored noise meeting reinforced Dalang’s condition. Recently, A. Agresti and M.C. Veraar [1, 2] demonstrated the existence, uniqueness, and regularity of solutions within critical spaces for stochastic reaction-diffusion equations with transport noise in periodic domains.
@@
    Non-local operators: Beyond the findings of B.-S. Han [27], research has predominantly focused on second-order elliptic operators. This observation prompts the question: Is it feasible to establish solvability for SRDAEs that involve infinitesimal generators of subordinate Brownian motions?
@@
[…]
@@ 1.3. Conditions on spatially homogeneous colored noise
This subsection details the conditions imposed on spatially homogeneous colored noise $W$ characterized by a nonnegative definite correlation $\pi$. Here is our main assumption on correlation $\pi$. Throughout this article, we always set
$\delta_{0}\in(0,1],\quad\gamma\in(0,1)\quad\text{and}\quad r\in[1,d/(d-\delta_{0}(1-\gamma))).$
@@
    The correlation $\pi$ satisfies that
    $\begin{cases}\int_{|x|<1}|x|^{2\alpha_{\delta_{0},\gamma,r}-d}\pi(\mathrm{d}x)<\infty,\quad&\text{if }0<\alpha_{\delta_{0},\gamma,r}<d/2,\\ \int_{|x|<1}\log(1/|x|)\pi(\mathrm{d}x)<\infty,\quad&\text{if }\alpha_{\delta_{0},\gamma,r}=d/2,\\ \text{no specific conditions on $\pi$,}\quad&\text{if }\alpha_{\delta_{0},\gamma,r}>d/2,\end{cases}$
@@
$\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}\ast R_{\delta_{0}(1-\gamma)})(x)\pi(\mathrm{d}x)=\int_{\mathbb{R}^{d}}R_{2\delta_{0}(1-\gamma)}(x)\pi(\mathrm{d}x)=:\nu_{\delta_{0},\gamma,1}<\infty,$
@@
The reinforced Dalang’s condition plays a crucial role in the study of the regularity of solutions to SPDEs driven by colored noise; see, for example, [18, 30, 55].
@@
Let $\delta_{0}\in(0,1]$, $\gamma\in(0,1)$, $r_{1},r_{2}\in[1,d/(d-\delta_{0}(1-\gamma)))$. Suppose that $\pi$ is a correlation satisfying Assumption 1.5 $(\delta_{0},\gamma,r_{2})$. If $r_{1}<r_{2}$, then $\pi$ also satisfies Assumption 1.5 $(\delta_{0},\gamma,r_{1})$.
@@
    for all $x\in\mathbb{R}^{d}\setminus\{0\}$ and $\beta\in(0,d)$,
    $R_{\beta}(x)\leq N(\beta,d)|x|^{\beta-d}1_{|x|\leq 2}+N(\beta,d)\mathrm{e}^{-\frac{|x|}{2}}1_{|x|>2}.$ (1.5)
@@
    for $|x|\leq 2$ and $\beta\in(0,d)$,
    $N(\beta,d)|x|^{\beta-d}\leq R_{\beta}(x).$ (1.6)
@@
$(4^{\prime})$ for $\kappa>0$ and $\beta\in(0,d)$, there exists $N=N(\beta,d,\kappa)>0$ such that for $|x|<\kappa$,
$N|x|^{\beta-d}\leq R_{\beta}(x).$ (1.7)
@@
Let $\delta_{0}\in(0,1]$, $\gamma\in(0,1)$, $r\in[1,d/(d-\delta_{0}(1-\gamma)))$ and $\pi$ be a tempered measure defined on $\mathbb{R}^{d}$. Then, the following conditions are equivalent:
@@
    The convolution $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})$ belongs to $L_{1}(\mathbb{R}^{d},\pi)$;
    $\nu_{\delta_{0},\gamma,r}:=\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)\pi(\mathrm{d}x)<\infty.$ (1.9)
@@
    The convolution $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})$ belongs to $L_{1}(B_{1}(0),\pi)$;
    $\int_{|x|<1}(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)\pi(\mathrm{d}x)<\infty.$ (1.10)
@@
The implication from condition $(ii)$ to $(iii)$ is straightforward. By examining the exponential decay of $R_{\delta_{0}(1-\gamma)}$ as $|x|\to\infty$, as outlined in (1.5) and corroborated by [9, Lemma 4.3], we infer that the convolution $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)$ exhibits similar exponential decay for large $x$. Furthermore, the characterization of $\pi$ as a tempered measure, detailed in [9, Definition 2.1], stipulates the existence of some $k\geq 0$ such that
$A_{\pi}:=\int_{\mathbb{R}^{d}}\frac{1}{(1+|x|^{2})^{k/2}}\pi(\mathrm{d}x)<\infty.$ (1.11)
This integral condition on $\pi$ robustly supports the assertion that $(iii)$ implies $(ii)$. We proceed by establishing the equivalence between conditions $(i)$ and $(iii)$ through three cases based on the value of $\alpha_{\delta_{0},\gamma,r}$.
@@
where the strict positivity and radial decrease of $R_{\delta_{0}(1-\gamma)}$ imply for $1/2<|y|<3$,
@@
$\nu_{\delta_{0},\gamma,r_{1}}$ $=\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r_{1}}\ast R_{\delta_{0}(1-\gamma)}^{r_{1}})(x)\pi(\mathrm{d}x)$ (1.14)
$\leq C_{0}\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\pi(\mathrm{d}x).$
Given that $(R_{\delta_{0}(1-\gamma)}^{r}\ast R_{\delta_{0}(1-\gamma)}^{r})(x)$ decays exponentially as $|x|\to\infty$ and considering $\pi$ as a tempered measure,
$\int_{\mathbb{R}^{d}}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\pi(\mathrm{d}x)<\infty\Longleftrightarrow\int_{|x|\leq 1}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\pi(\mathrm{d}x)<\infty.$ (1.15)
@@
$\int_{|x|\leq 1}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)^{\frac{r_{1}^{2}}{r_{2}^{2}}}\pi(\mathrm{d}x)$ (1.16)
$\leq A_{\pi}^{1-\frac{r_{1}^{2}}{r_{2}^{2}}}\left(\int_{|x|\leq 1}(R_{\delta_{0}(1-\gamma)}^{r_{2}}\ast R_{\delta_{0}(1-\gamma)}^{r_{2}})(x)(1+|x|^{2})^{\frac{k}{2}\left(\frac{r_{2}^{2}}{r_{1}^{2}}-1\right)}\pi(\mathrm{d}x)\right)^{\frac{r_{1}^{2}}{r_{2}^{2}}}$
$\leq N(A_{\pi},k,r_{1},r_{2})\nu_{\delta_{,}\gamma,r_{2}}^{\frac{r_{1}^{2}}{r_{2}^{2}}},$
where $A_{\pi}$ is the constant defined in (1.11). Consequently,